# Maximum Quantum Work at Criticality: Stirling Engines and Fibonacci-Lucas Degeneracies

### v1

@@ II Model
$U_{\lambda_{\text{crit}}}(T\rightarrow 0)=\epsilon_{\text{ground}}^{\text{(crit)}}$ $\implies$ $S_{\lambda_{\text{crit}}}(T\rightarrow 0)=k_{B}\ln\quantity(g_{\text{crit}})$ (6)
$U_{\lambda_{H}}(T\rightarrow 0)=\epsilon_{0}$ $\implies$ $S_{\lambda_{H}}(T\rightarrow 0)=k_{B}\ln\quantity(g_{0})$
@@ III.2 Quantum Stirling Engines at Quantum Criticality
    $U_{\text{crit}}^{(B)}(T_{H})U_{\lambda_{\text{crit}}}^{(C)}(T_{L})$ $=\epsilon_{\text{ground}}^{\text{(crit)}},$ $S_{\lambda_{\text{crit}}}^{(C)}(T_{L})=k_{B}\ln\quantity(g_{\text{crit}}).$ (16)
@@
$Q_{AB}$ $=T_{H}\quantity[S_{\lambda_{\text{crit}}}^{(B)}(T_{H})-S_{\lambda_{H}}^{(A)}(T_{H})]=k_{B}T_{H}[\ln(g_{\text{crit}})-\ln(g_{0})]=k_{B}T_{H}\ln\quantity(\frac{g_{\text{crit}}}{g_{0}}),$ (18)
$Q_{CD}$ $=T_{L}\quantity[S_{\lambda_{H}}^{(D)}(T_{L})-S_{\lambda_{\text{crit}}}^{(C)}(T_{L})]=k_{B}T_{L}[\ln(g_{0})-\ln(g_{\text{crit}})]=-k_{B}T_{L}\ln\quantity(\frac{g_{\text{crit}}}{g_{0}}),$ (20)
$Q_{DA}$ $=U_{\lambda_{H}}^{(A)}(T_{H})-U_{\lambda_{H}}^{(D)}(T_{L})=0.$ (21)
@@
$\eta_{C}=\frac{\displaystyle{k_{B}T_{H}\ln\left(\frac{g_{\text{crit }}}{g_{0}}\right)}-k_{B}T_{L}\ln\left(\frac{g_{\text{crit }}}{g_{0}}\right)}{\displaystyle k_{B}T_{H}\ln\left(\frac{g_{\text{crit }}}{g_{0}}\right)}=1-\frac{T_{L}}{T_{H}}=\frac{1}{1+\frac{TL}{\delta}}.$ (24)
@@ III.3 Thermal Excitations and Small-Gap Effects
$F_{\lambda_{H}}(T)$ $=-k_{B}T\ln\left(Z_{\lambda_{H}}(T)\right)=-k_{B}T\ln\left(e^{-\frac{\displaystyle\epsilon_{0}}{k_{B}T}}\left[g_{0}+g_{1}e^{-\frac{\Delta}{k_{B}T}}\right]\right),$ (32)
@@
$Q_{AB}$ $=k_{B}T_{H}\ln\left(g_{\text{crit}}\right)-T_{H}S_{\lambda_{H}}(T_{H}),$ $Q_{AD}=U_{\lambda_{H}}(T_{H})-\epsilon_{0}.$ (35)
@@
$Q_{AB}$ $=T_{H}\left[S_{\lambda_{\text{cri t}}}^{(B)}(T_{H})-S_{\lambda_{H}}^{(A)}(T_{H})\right]=k_{B}T_{H}\ln\left(g_{\text{crit}}\right)-U_{\lambda_{H}}^{(A)}(T_{H})+F^{(A)}_{\lambda_{H}}(T_{H}),$ (36)
@@
$Q_{CD}$ $=T_{L}\left[S_{\lambda_{H}}^{(D)}(T_{L})-S_{\lambda_{\text{crit}}}^{(C)}(T_{L})\right]=k_{B}T_{L}\left[\ln(g_{0})-\ln(g_{\text{crit}})\right]=-k_{B}T_{L}\ln\left(\frac{g_{\text{crit}}}{g_{0}}\right),$ (38)
@@
IV.1 Validation Across Reported Quantum Systems
IV.1.1 Two spin-1/2 systems
$\hat{\mathcal{H}}_{2\,spin-1/2}$ $=J(\hat{S}_{1}^{x}\hat{S}_{2}^{x}+\hat{S}_{1}^{y}\hat{S}_{2}^{y})+J_{z}\hat{S}_{1}^{z}\hat{S}_{2}^{z}+B_{z}\hat{S}_{\text{tot }}^{z}+K_{y}(\hat{S}_{\text{tot }}^{y})^{2}+D\left(\hat{\mathbf{S}}_{1}\times\hat{\mathbf{S}}_{2}\right)\cdot\hat{\mathbf{z}}$. (44)
@@ IV.1.1 Two spin-1/2 systems
IV.1.2 Nth spin-s system
$\hat{\mathcal{H}}_{\mathrm{spin-s}}^{(N)}$ $=J\sum_{i=1}^{N-1}\hat{\mathbf{S}}_{i}\cdot\hat{\mathbf{S}}_{i+1}+B\hat{S}_{\mathrm{tot}}^{z}+K_{y}\left(\hat{S}_{\mathrm{tot}}^{y}\right)^{2}+K_{z}\sum_{i=1}^{N}\left(\hat{S}_{i}^{z}\right)^{2}.$ (45)
@@ IV.2 Quantum Ising model at criticality with Fibonacci-Lucas Degeneracies
Figure 4: Quantum Stirling work $W$ [meV] as a function of the number of sites NN for the antiferromagnetic Ising chain and ring ($J=1$), evaluated at the QCP $\lambda_{L}=B/J=1$ and a high-field parameter $\lambda_{H}=B/J=3$, with reservoir temperatures $T_{L}=0.1$ and $T_{H}=0.2$. In all cases the system operates at Carnot efficiency $\eta_{C}=0.5$. Red (black) dots represent simulated cycles for the chain (ring) system. Yellow and green markers show the predictions of Eqs. (49-50) based on Fibonacci and Lucas degeneracies at the QCP, respectively, while blue (orange) lines correspond to the large-NN asymptotic expressions in Eqs. (51-52).
@@
Eqs. (51-52) show that both cases scale linearly with $N$, but the chain includes a constant offset of $\ln\sqrt{5}$, highlighting a residual distinction between quantum thermodynamics and its classical counterpart. However, in the thermodynamic limit $N\ggg 1$, the constant term becomes negligible, and both chain and ring topologies recover classical extensivity, yielding the same work output determined solely by the system size. This behavior is consistent with results reported in the literature for the classical Ising model.
@@ Acknowledgements.
Author Contributions
Bastian Castorene  — Conceptualization (lead), Data curation (lead), Formal analysis (lead), Investigation (lead), Methodology (lead), Visualization (lead), Writing – original draft (lead), Writing – review & editing (lead).
Martin HvE Groves  — Formal analysis (equal), Investigation (supporting), Validation (equal), Writing – review & editing (lead).
Francisco J. Peña  — Investigation (supporting), Supervision (equal), Validation (lead), Writing – review & editing (lead).
Eugenio E. Vogel  — Conceptualization (supporting), Funding acquisition (lead), Supervision (equal), Validation (equal), Writing – review & editing (supporting).
Patricio Vargas  — Conceptualization (lead), Methodology (lead), Supervision (lead), Writing – original draft (supporting), Writing – review & editing (equal).
Data Availability Statement
Data sharing is not applicable to this article as no new data were created or analyzed in this study.

### v3

@@ II Model
$U_{\lambda_{\text{crit}}}(T\rightarrow 0)=\epsilon_{\text{ground}}^{\text{(crit)}}\implies S_{\lambda_{\text{crit}}}(T\rightarrow 0)=k_{B}\ln\quantity(g_{\text{crit}}),$ (6)
$U_{\lambda_{H}}(T\rightarrow 0)=\epsilon_{0}\implies S_{\lambda_{H}}(T\rightarrow 0)=k_{B}\ln\quantity(g_{0}).$
@@ III.2 Quantum Stirling Engines at Quantum Criticality
    $U_{\lambda_{\text{crit}}}^{(C)}(T_{L})$ $=\epsilon_{\text{ground}}^{\text{(crit)}},$ $S_{\lambda_{\text{crit}}}^{(C)}(T_{L})=k_{B}\ln\quantity(g_{\text{crit}}).$ (16)
@@
$\begin{aligned} Q_{AB}&=T_{H}\quantity[S_{\lambda_{\text{crit}}}^{(B)}(T_{H})-S_{\lambda_{H}}^{(A)}(T_{H})]\\ &=k_{B}T_{H}\ln\quantity(\frac{g_{\text{crit}}}{g_{0}}),\end{aligned}$ (18)
$\begin{aligned} Q_{BC}&=U_{\lambda_{\text{crit}}}^{(C)}(T_{L})-U_{\lambda_{\text{crit}}}^{(B)}(T_{H})=0,\end{aligned}$ (19)
$\begin{aligned} Q_{CD}&=T_{L}\quantity[S_{\lambda_{H}}^{(D)}(T_{L})-S_{\lambda_{\text{crit}}}^{(C)}(T_{L})]\\ &=-k_{B}T_{L}\ln\quantity(\frac{g_{\text{crit}}}{g_{0}}),\end{aligned}$ (20)
$\begin{aligned} Q_{DA}&=U_{\lambda_{H}}^{(A)}(T_{H})-U_{\lambda_{H}}^{(D)}(T_{L})=0.\end{aligned}$ (21)
@@
$\eta_{C}$ $=\frac{\displaystyle{k_{B}T_{H}\ln\left(\frac{g_{\text{crit }}}{g_{0}}\right)}-k_{B}T_{L}\ln\left(\frac{g_{\text{crit }}}{g_{0}}\right)}{\displaystyle k_{B}T_{H}\ln\left(\frac{g_{\text{crit }}}{g_{0}}\right)}$ (24)
$=1-\frac{T_{L}}{T_{H}}=\frac{1}{\displaystyle 1+\frac{T_{L}}{\delta}}.$ (25)
@@ III.3 Thermal Excitations and Small-Gap Effects
$Q_{AB}$ $=k_{B}T_{H}\ln\left(g_{\text{crit}}\right)-T_{H}S_{\lambda_{H}}(T_{H}),$ (36)
$Q_{AD}$ $=U_{\lambda_{H}}(T_{H})-\epsilon_{0}.$ (37)
@@
$\begin{aligned} Q_{AB}&=T_{H}\left[S_{\lambda_{\text{cri t}}}^{(B)}(T_{H})-S_{\lambda_{H}}^{(A)}(T_{H})\right]\\ &=k_{B}T_{H}\ln\left(g_{\text{crit}}\right)-U_{\lambda_{H}}^{(A)}(T_{H})+F^{(A)}_{\lambda_{H}}(T_{H}),\end{aligned}$ (38)
@@
$\begin{aligned} Q_{CD}&=T_{L}\left[S_{\lambda_{H}}^{(D)}(T_{L})-S_{\lambda_{\text{crit}}}^{(C)}(T_{L})\right]\\ &=-k_{B}T_{L}\ln\left(\frac{g_{\text{crit}}}{g_{0}}\right),\end{aligned}$ (40)
@@
IV.1 Validation Across Reported Quantum Systems
IV.1.1 Two spin-1/2 systems
$\begin{aligned} \hat{\mathcal{H}}_{2\,spin-1/2}=&J(\hat{S}_{1}^{x}\hat{S}_{2}^{x}+\hat{S}_{1}^{y}\hat{S}_{2}^{y})+J_{z}\hat{S}_{1}^{z}\hat{S}_{2}^{z}+B_{z}\hat{S}_{\text{tot }}^{z}+\\ &+K_{y}(\hat{S}_{\text{tot }}^{y})^{2}+D\left(\hat{\mathbf{S}}_{1}\times\hat{\mathbf{S}}_{2}\right)\cdot\hat{\mathbf{z}}\end{aligned}$. (46)
IV.1.2 Nth spin-s system
$\begin{aligned} \hat{\mathcal{H}}_{\mathrm{spin-s}}^{(N)}=&J\sum_{i=1}^{N-1}\hat{\mathbf{S}}_{i}\cdot\hat{\mathbf{S}}_{i+1}+B\hat{S}_{\mathrm{tot}}^{z}+\\ &+K_{y}\left(\hat{S}_{\mathrm{tot}}^{y}\right)^{2}+K_{z}\sum_{i=1}^{N}\left(\hat{S}_{i}^{z}\right)^{2}\end{aligned}.$ (47)
@@ IV.2 Quantum Ising model at criticality with Fibonacci-Lucas Degeneracies
Figure 4: Quantum Stirling work $W$ [meV] as a function of the number of sites NN for the antiferromagnetic Ising chain and ring ($J=1$), evaluated at the QCP $\lambda_{L}=B/J=1$ and a high-field parameter $\lambda_{H}=B/J=3$, with reservoir temperatures $T_{L}=0.1$ and $T_{H}=0.2$. In all cases the system operates at Carnot efficiency $\eta_{C}=0.5$. Red (black) dots represent simulated cycles for the chain (ring) system. Yellow and green markers show the predictions of Eqs. (51-52) based on Fibonacci and Lucas degeneracies at the QCP, respectively, while blue (orange) lines correspond to the large-NN asymptotic expressions in Eqs. (53-54).
@@
Eqs. (53-54) show that both cases scale linearly with $N$, but the chain includes a constant offset of $\ln\sqrt{5}$, highlighting a residual distinction between quantum thermodynamics and its classical counterpart. However, in the thermodynamic limit $N\ggg 1$, the constant term becomes negligible, and both chain and ring topologies recover classical extensivity, yielding the same work output determined solely by the system size. This behavior is consistent with results reported in the literature for the classical Ising model.